OpenStudy (anonymous):

What cosine function represents an amplitude of 2, a period of 2π, a horizontal shift of π, and a vertical shift of −1? f(x) = −1 cos πx + 2 f(x) = −1 cos (x − π) + 2 f(x) = 2 cos (x − π) − 1 f(x) = 2 cos πx − 1

OpenStudy (anonymous):

WILL GIVE MEDAL TO WHOEVER IS CORRECT!

hartnn (hartnn):

general form of cos function : \(\Large f(x) = A \cos (\dfrac{2\pi }{T}x- \alpha )+b\) where A = Amplitude T =period \(\alpha\) = horizontal shift b = vertical shift just plug in values!

OpenStudy (anonymous):

thank you! @hartnn

hartnn (hartnn):

welcome ^_^